# A framework for the valuation of insurance liabilities by production cost

### v1

@@ 6 Valuation of insurance liabilities under Solvency II and the SST
In the following we sketch a way to derive the valuation of insurance liabilities in Solvency II ([3], [6]) and the SST ([9]) from the framework, in principle by specifying the applicable fulfillment and financiability condition. However, to get to the usual representation in Solvency II and the SST, we use additional simplifying assumptions, especially on the production strategies, as we show.
@@
In solvency systems such as Solvency II and the SST, an insurance company is solvent at a reference date $i=0$ if the available capital $AC_{0}$ (Eligible Own Funds in Solvency II, Risk-Bearing Capital in the SST) is at least as large as the regulatory required solvency capital $SCR_{0}$ (Solvency Capital Requirement in Solvency II, Target Capital in the SST). The available capital $AC_{0}$ is roughly the value of the assets minus the value of the liabilities. The required capital $SCR_{0}$ is broadly derived from the discounted one-year change in available capital $AC_{i}$ from date 00 to $1$:
$SCR_{0}=\rho[(1+r_{0,1})^{-1}\cdot AC_{1}-AC_{0}]$ (24)
@@
In Solvency II and the SST, the value of the insurance liabilities $\mathcal{L}$ at date $i=0$ is given by the sum of best estimate $BEL_{0}$ and risk margin $RM_{0}$ (market value margin in the SST). The best estimate $BEL_{0}$ is roughly defined as the expected value of all discounted cash flows for producing the liabilities with the exception of the capital cost (and is typically calculated assuming perfect production). The risk margin $RM_{0}$ accounts for the capital cost of the required capital $SCR_{i}$ over the lifetime $i\in\{0,\ldots,T\}$ of the insurance liabilities under specific assumptions.
To derive this from the framework, we assume for simplicity that there are no cash flows for dates $t\in D$ with $i<t<i+1$. We assume that, for any $i\in\{0,\ldots,T-1\}$, there exists a “risk-free one-year zero-coupon bond” with rate $r_{i,i+1}$ from date $i$ to $i+1$ in the selected currency, i.e. a tradable $k\equiv k_{i}$ with $S_{i}^{k}=(1+r_{i,i+1})^{-1}$, $S_{i+1}^{k}=0$, and $\tilde{Z}_{i+1}^{k}=Z_{i+1}^{k}=1$. We denote by $AC_{i}^{pre}$ and $AC_{i}^{post}$ the available capital before and after liability cash flows $X_{i}^{\mathcal{L}}$, capital payback, and capital raise at date $i$, respectively, using Definition 4.1. So, for a production strategy $\phi$ with illiquid assets $\psi$, we have $AC_{i}^{pre}=A_{i}^{\prime}-L_{i}=\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}-X_{i}^{\mathcal{L}}-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and $AC_{i}^{post}=AC_{i}^{pre}-(A_{i}^{\prime}-L_{i})_{+}+C_{i}=C_{i}$ in case $A_{i}^{\prime}\geq L_{i}$. The value $A_{i}=\phi_{\gamma(i)}\cdot S_{i}$ of the assets in the “post”-case is the sum of the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and the capital $C_{i}$ (Section 4.1). We define the fulfillment condition by $\rho[AC_{1}^{pre}]\leq 0$ for a risk measure $\rho$ that is translation invariant ($\rho[Y+a]=\rho[Y]-a$ for any $a\in\mathbb{R}$) and positively homogeneous ($\rho[a\cdot Y]=a\cdot\rho[Y]$ for any $a\geq 0$). The risk measures in Solvency II and the SST satisfy these conditions. We define the financiability condition by the requirement that the capital payoff $C_{1}^{\prime}=(AC_{1}^{pre})_{+}$ provide at least an expected excess return $\eta\geq 0$ over risk free on the capital $C_{0}$, i.e. $\mathbb{E}[(AC_{1}^{pre})_{+}]\geq(1+r_{0,1}+\eta)\cdot C_{0}$. This satisfies the two conditions from Definition 4.3 and is positively homogeneous, but not necessarily consistent with and neutral to the tradables (and is not necessarily a realistic condition). We write $\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]$ with $M_{1}=\{A_{1}^{\prime}\geq L_{1}\}$.
@@
Assuming it exists, let $SCR_{0}$ be the capital $C_{0}=AC_{0}^{post}$ for which the fulfillment condition is satisfied with equality, i.e. $\rho[AC_{1}^{pre}]=0$. We can multiply the latter expression by $(1+r_{0,1})^{-1}$, add $SCR_{0}=AC_{0}^{post}$, and use that $\rho$ is translation invariant and positively homogeneous to get
$SCR_{0}=\rho[(1+r_{0,1})^{-1}\cdot AC_{1}^{pre}-AC_{0}^{post}]$ (25)
So the fulfillment condition can be seen as a special case of formula (24) for the regulatory required solvency capital in Solvency II and the SST. It can of course not directly be used to calculate $SCR_{0}$ as $AC_{0}^{post}=SCR_{0}$ appears on the right-hand side.
In line with the recursively backward-in-time approach in the framework, we assume that we are given the liability value $L_{1}=X_{1}^{\mathcal{L}}+\bar{v}_{1}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=X_{1}^{\mathcal{L}}+BEL_{1}+RM_{1}$ with best estimate $BEL_{1}$ and risk margin $RM_{1}$ at date $i=1$. The objective is to derive a production strategy $\phi$ from date $i=0$ to $i=1$ and the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=BEL_{0}+RM_{0}$ at date $i=0$. In our framework, $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ is calculated in two steps, without calculating $BEL_{0}$ and $RM_{0}$ separately. The first step is satisfying the fulfillment condition $\rho[AC_{1}^{pre}]\leq 0$, in particular determining the value $A_{0}=\phi_{\gamma(0)}\cdot S_{0}$ of the total assets required at $i=0$ to satisfy the fulfillment condition. In the special case that the assets are only invested in the risk-free one-year zero-coupon bond, i.e. $A_{1}^{\prime}=(1+r_{0,1})\cdot A_{0}$, the value $A_{0}=(1+r_{0,1})^{-1}\cdot\rho(-L_{1})$ solves the fulfillment condition with equality, $\rho(A_{1}^{\prime}-L_{1})=0$, as $\rho$ is translation-invariant, and we also get $M_{1}=\{L_{1}\leq\rho[-L_{1}]\}$. The second step is solving for the capital $C_{0}=SCR_{0}$ in the financiability condition $\mathbb{E}[(AC_{1}^{pre})_{+}]=(1+r_{0,1}+\eta)\cdot C_{0}$ with equality, from which we get the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=A_{0}-SCR_{0}$ at date $i=0$. In the above special case, $(1+r_{0,1}+\eta)\cdot SCR_{0}=\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]=P[L_{1}\leq\rho[-L_{1}]]\cdot(1+r_{0,1})\cdot A_{0}-\mathbb{E}[1_{\{L_{1}\leq\rho[-L_{1}]\}}\cdot L_{1}]$. Inserting the resulting expression for $SCR_{0}$ and the above expression for $A_{0}$ into $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=A_{0}-SCR_{0}$, we get an explicit formula for the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in terms of the liabilities $L_{1}$ (see Theorem 4 of [8]).
For the usual presentation in Solvency II and the SST, we need to determine the two components $BEL_{0}$ and $RM_{0}$ of the production cost, where $RM_{0}$ should correspond to the capital cost for the required capital $SCR_{i}$ for all $i=0,1,\dots,T-1$. To this end, we make the following two additional simplifying assumptions. First, we assume that the assets corresponding to each summand in $A_{0}=SCR_{0}+BEL_{0}+RM_{0}$ are invested separately, with values at date $1$ denoted by $A_{1}^{\prime}=A_{1}^{SCR}+A_{1}^{BEL}+A_{1}^{RM}$. Second, the assets corresponding to $SCR_{0}$ and $RM_{0}$ are invested in the risk-free zero-coupon bond, so $A_{1}^{SCR}=(1+r_{0,1})\cdot SCR_{0}$ and $A_{1}^{RM}=(1+r_{0,1})\cdot RM_{0}$. Then, we get for the fulfillment condition with equality, using that $\rho$ is translation-invariant, $0=\rho[AC_{1}^{pre}]=-(1+r_{0,1})\cdot SCR_{0}+\rho[A_{1}^{BEL}+A_{1}^{RM}-L_{1}]$. We find from this that the required capital $SCR_{0}=(1+r_{0,1})^{-1}\cdot\rho[A_{1}^{BEL}+A_{1}^{RM}-L_{1}]$ arises from the mismatch between the value $A_{1}^{BEL}+A_{1}^{RM}$ of the “replicating portfolio” and the value $L_{1}$ of the liabilities. This formula cannot directly be used to calculate $SCR_{0}$, but we further get $SCR_{0}=-RM_{0}+(1+r_{0,1})^{-1}\cdot\rho[A_{1}^{BEL}-L_{1}]$ and $M_{1}=\{A_{1}^{BEL}-L_{1}\geq-\rho[A_{1}^{BEL}-L_{1}]\}$.
For the financiability condition, we have $\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]$ with $A_{1}^{\prime}=(1+r_{0,1})\cdot(SCR_{0}+RM_{0})+A_{1}^{BEL}$ and $L_{1}=X_{1}^{\mathcal{L}}+BEL_{1}+RM_{1}$. The split into $BEL_{0}$ and $RM_{0}$ needs to be defined by a suitable “separation condition”. The best estimate is intended to account for all cash flows other than capital cost, so we assume that the condition is $\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{BEL}-X_{1}^{\mathcal{L}}-BEL_{1})]=0$ (which, however, involves $RM_{1}$ in $M_{1}$). With this, the financiability condition becomes
@@
Solving this for $RM_{0}$, the risk margin $RM_{0}$ can be written as the sum of a term that could be interpreted as capital cost on the required capital $SCR_{0}$ for the period from 00 to $1$ and a term containing $RM_{1}$, i.e. the capital cost for subsequent periods. To actually derive $BEL_{0}$ and $RM_{0}$, the first step is determining the best estimate $BEL_{0}$ such that the separation condition holds, which does not involve $SCR_{0}$ and $RM_{0}$. Then, the risk margin $RM_{0}$ (and if required $SCR_{0}$) can be calculated from equation (26) and the equation $SCR_{0}=-RM_{0}+(1+r_{0,1})^{-1}\cdot\rho[A_{1}^{BEL}-L_{1}]$ from the fulfillment condition above.
To get closer to the usual presentation of Solvency II and SST valuation, we make a further simplification: the financiability condition with equality is replaced by the stricter condition $\mathbb{E}[AC_{1}^{pre}]=(1+r_{0,1}+\eta)\cdot SCR_{0}$ by removing $()_{+}$. This gives a lower required capital $SCR_{0}$ and thus a higher production cost than without the simplification and in this sense an upper bound. Formula (26) simplifies to give for the risk margin $RM_{0}$:
@@
So the risk margin is the sum of the discounted capital cost for the period from 00 to $1$ and the discounted expected capital cost for the periods after $1$, which is close to corresponding expressions for Solvency II (Article 37 of [6]) and the SST. As above, $RM_{0}$ cannot directly be calculated from this even after $BEL_{0}$ has been determined because it appears also in $SCR_{0}$.
With the simplified stricter financiability condition and further assuming that also the assets for the best estimate are invested in the risk-free one-year zero-coupon bond, i.e. $A_{1}^{BEL}=(1+r_{0,1})\cdot BEL_{0}$, the correspondingly simplified separation condition $\mathbb{E}[A_{1}^{BEL}-X_{1}^{\mathcal{L}}-BEL_{1}]=0$ directly gives

### v2

@@ 5 Valuation of insurance liabilities under Solvency II and in the SST
In the following we set out a way to derive a valuation of insurance liabilities corresponding to Solvency II ([3], [6]) and the SST ([9]) from the framework. This involves specifying the applicable fulfillment and financiability condition and production strategies, as well as additional simplifying assumptions.
@@
Solvency II and SST. We recall that an insurance company is solvent at a reference date $i=0$ if the available capital $AC_{0}$ (Eligible Own Funds in Solvency II, Risk-Bearing Capital in the SST) is at least as large as the regulatory required solvency capital $SCR_{0}$ (Solvency Capital Requirement in Solvency II, Target Capital in the SST):
$AC_{0}\geq SCR_{0}$ (30)
The available capital $AC_{0}$ is roughly the value of the assets minus the value of the liabilities. The required capital $SCR_{0}$ is largely derived from the discounted one-year change in available capital:
$SCR_{0}=\rho((1+r_{0,1})^{-1}\cdot AC_{1}-AC_{0})$ (31)
@@
The value of the insurance liabilities $\mathcal{L}$ at date $i=0$ is given by the sum of best estimate $BEL_{0}$ and risk margin $RM_{0}$ (market value margin in the SST):
$BEL_{0}+RM_{0}$ (32)
@@
The best estimate $BEL_{0}$ is roughly defined as the expected value of all discounted cash flows needed for producing the liabilities with the exception of capital cost (and typically without considering fulfillment in only “sufficiently many cases”). The risk margin $RM_{0}$ accounts for the capital cost of the required capital $SCR_{i}$ over the lifetime $i\in\{0,\ldots,T\}$ of the insurance liabilities, specified through a cost-of-capital rate $\eta>0$ in excess of the risk-free rate. In Solvency II regulation, the following formula for the risk margin is provided, with $\eta=CoC$ (Article 37 of [6]):
$RM_{0}=CoC\cdot\sum_{i\geq 0}\frac{SCR_{i}}{(1+r_{0,i+1})^{i+1}}$ (33)
The $SCR_{i}$ are calculated under specific assumptions including no new business, available capital $AC_{i}$ equal to required capital $SCR_{i}$, and an investment strategy that is adapted to the insurance liabilities, e.g. under Solvency II so that market risk is “minimized”.
Integration into the framework. For the fulfillment condition, we get from (30) and (31) using translation invariance and positive homogeneity of $\rho$ that $AC_{0}\geq(1+r_{0,1})^{-1}\cdot\rho(AC_{1})+AC_{0}$, which suggests that the fulfillment condition is $\rho(AC_{1})\leq 0$. To be more precise, in our setting of production strategies, $AC_{1}$ here and in (31) is, for $i=1$, the available capital $AC_{i}^{pre}=A_{i}^{\prime}-L_{i}$ before liability cash flows $X_{i}^{\mathcal{L}}$, capital payback, and capital raise at date $i$, respectively, using Definition 3.1. In contrast, $AC_{0}$ above and in (31) is, for $i=0$, the available capital $AC_{i}^{post}=AC_{i}^{pre}-(A_{i}^{\prime}-L_{i})_{+}+C_{i}=C_{i}$ after these steps, in case $A_{i}^{\prime}\geq L_{i}$. The value $\phi_{\gamma(i)}\cdot S_{i}$ of the assets in the “post”-case is the sum of the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and the capital $C_{i}$.
For the financiability condition, we interpret the cost-of-capital rate $\eta>0$ as the minimum (real-world) expected return on the required capital $SCR_{0}$ in excess of the risk-free return. We thus assume:
  1. (a)
    The fulfillment condition is
    $\rho(AC_{1}^{pre})\leq 0$ (34)
    with a translation-invariant and positively homogeneous risk measure $\rho$.
  2. (b)
    The financiability condition is the requirement that the capital payoff $C_{1}^{\prime}=(AC_{1}^{pre})_{+}$ provide at least a (real-world) expected excess return $\eta\geq 0$ over risk free on the capital $C_{0}$, i.e.
    $\mathbb{E}[(AC_{1}^{pre})_{+}]\geq(1+r_{0,1}+\eta)\cdot C_{0}$ (35)
    We write $\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]$ with $M_{1}=\{A_{1}^{\prime}\geq L_{1}\}$.
The financiability condition (35) satisfies the two conditions from Definition 3.6 and is positively homogeneous but not necessarily consistent with and neutral to the tradables (and may not be a realistic condition).
We further assume that there are no cash flows for dates $t\in D$ with $i<t<i+1$ and that, for any $i\in\{0,\ldots,T-1\}$, there exists a “risk-free one-year zero-coupon bond” with rate $R_{i,i+1}$ from date $i$ to $i+1$ in the selected currency, i.e. a tradable $k\equiv k_{i}$ with $S_{i}^{k}=(1+R_{i,i+1})^{-1}$, $S_{i+1}^{k}=0$, and $\tilde{Z}_{i+1}^{k}=Z_{i+1}^{k}=1$. We have $r_{0,1}=R_{0,1}$.
@@
To investigate valuation in the framework under the above assumptions, we use the recursively backward-in-time approach of the framework. We consider three stages of successively more simplifications, with each stage moving closer to the usual presentation of Solvency II and SST valuation. In the first stage, the value is calculated as a whole, with no separate calculation of $BEL_{0}$ and $RM_{0}$. In the second stage, $BEL_{0}$ and $RM_{0}$ are calculated separately, based on restricting the production strategies. We note that the separate calculation creates additional work and would not be required within the framework. In the third stage, the financiability condition (35) is replaced by a simpler condition.
To simplify, we assume that we are given the liability value $L_{1}=X_{1}^{\mathcal{L}}+\bar{v}_{1}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=X_{1}^{\mathcal{L}}+BEL_{1}+RM_{1}$ with best estimate $BEL_{1}$ and risk margin $RM_{1}$ at date $i=1$. The objective is to derive the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=BEL_{0}+RM_{0}$ at date $i=0$.
First stage. Following the framework, $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ is calculated in the two steps set out in the introduction:
  1. 1:
    Determine the value $A_{0}=\phi_{\gamma(0)}\cdot S_{0}$ of the total assets required at date $i=0$ and a production strategy $\phi$ from $i=0$ to $i=1$ such that the fulfillment condition $\rho(AC_{1}^{pre})\leq 0$ is satisfied at $i=1$.
In the special case that the assets are only invested in the risk-free one-year zero-coupon bond, i.e. $A_{1}^{\prime}=(1+r_{0,1})\cdot A_{0}$, the value $A_{0}=(1+r_{0,1})^{-1}\cdot\rho(-L_{1})$ solves the fulfillment condition with equality, $\rho(A_{1}^{\prime}-L_{1})=0$, as $\rho$ is translation-invariant. We also get $M_{1}=\{L_{1}\leq\rho(-L_{1})\}$.
The second step is:
  1. 2:
    Solve for the capital $C_{0}=SCR_{0}$ in the financiability condition $(1+r_{0,1}+\eta)\cdot C_{0}=\mathbb{E}[(AC_{1}^{pre})_{+}]=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]$ with equality. From this, we get the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=A_{0}-SCR_{0}$ at date $i=0$.
In the above special case, we have $\mathbb{E}[1_{M_{1}}\cdot A_{1}^{\prime}]=(1+r_{0,1})\cdot A_{0}\cdot\mathbb{E}[1_{M_{1}}]=(1+r_{0,1})\cdot A_{0}\cdot P[L_{1}\leq\rho(-L_{1})]$. Hence, for the financiability condition, $(1+r_{0,1}+\eta)\cdot SCR_{0}=\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{\prime}-L_{1})]=(1+r_{0,1})\cdot A_{0}\cdot P[L_{1}\leq\rho(-L_{1})]-\mathbb{E}[1_{\{L_{1}\leq\rho(-L_{1})\}}\cdot L_{1}]$. Inserting the resulting expression for $SCR_{0}$ and the above expression $A_{0}=(1+r_{0,1})^{-1}\cdot\rho(-L_{1})$ into $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=A_{0}-SCR_{0}$, we get an explicit formula for the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ in terms of $L_{1}$ (see Theorem 4 of [8]).
Second stage. We calculate the two components $BEL_{0}$ and $RM_{0}$ of the production cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$, where $RM_{0}$ should correspond to the capital cost. For this, we make the following two simplifying assumptions:
  1. (c)
    The assets corresponding to each summand in $A_{0}=BEL_{0}+RM_{0}+SCR_{0}$ are invested separately, with the values at date $1$ denoted by $A_{1}^{\prime}=A_{1}^{BEL}+A_{1}^{RM}+A_{1}^{SCR}$.
  2. (d)
    The assets corresponding to $SCR_{0}$ and $RM_{0}$ are invested in the risk-free zero-coupon bond, so $A_{1}^{SCR}=(1+r_{0,1})\cdot SCR_{0}$ and $A_{1}^{RM}=(1+r_{0,1})\cdot RM_{0}$.
Inserting $A_{1}^{SCR}=(1+r_{0,1})\cdot SCR_{0}$ into the fulfillment condition with equality $\rho(AC_{1}^{pre})=0$, using translation-invariance and solving for $SCR_{0}$, we find that the required capital $SCR_{0}=(1+r_{0,1})^{-1}\cdot\rho(A_{1}^{BEL}+A_{1}^{RM}-L_{1})$ arises from the mismatch between the value $A_{1}^{BEL}+A_{1}^{RM}$ of the “replicating portfolio” and the value $L_{1}$ of the liabilities. However, this formula cannot directly be used to calculate $SCR_{0}$. Further inserting $A_{1}^{RM}=(1+r_{0,1})\cdot RM_{0}$ and using translation invariance:
$(1+r_{0,1})\cdot(RM_{0}+SCR_{0})=\rho(A_{1}^{BEL}-L_{1})$ (36)
Inserting this into $A_{1}^{\prime}=A_{1}^{BEL}+(1+r_{0,1})\cdot(RM_{0}+SCR_{0})$, we get $M_{1}=\{A_{1}^{\prime}\geq L_{1}\}=\{A_{1}^{BEL}-L_{1}\geq-\rho(A_{1}^{BEL}-L_{1})\}$.
The financiability condition with equality is that $(1+r_{0,1}+\eta)\cdot SCR_{0}$ needs to be equal to
$\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{BEL}+A_{1}^{RM}+A_{1}^{SCR}-X_{1}^{\mathcal{L}}-BEL_{1}-RM_{1})]$ (37)
To split the value into $BEL_{0}$ and $RM_{0}$, considering that the best estimate is intended to account for the expected cash flows other than capital cost in the “sufficiently many cases”, we require $A_{1}^{BEL}$ to satisfy:
  1. (e)
    The best estimate is defined by the condition: [^8] This condition involves the risk margin $RM_{1}$ through $M_{1}$, so it is not totally independent of the capital cost.
    $\mathbb{E}[1_{M_{1}}\cdot(A_{1}^{BEL}-X_{1}^{\mathcal{L}}-BEL_{1})]=0$ (38)
This can be solved for $BEL_{0}$ and its investment strategy from date 00 to $1$ without knowing $SCR_{0}$ or $RM_{0}$ because of the expression for $M_{1}$ derived above.
For calculating the risk margin $RM_{0}$ (and if required $SCR_{0}$), we insert (38) into (37) and derive from the financiability condition that:
@@
Solving this for the risk margin $RM_{0}$, it can be written as the sum of a term that can be interpreted as capital cost on the required capital $SCR_{0}$ for the period from 00 to $1$ and a term containing $RM_{1}$, i.e. the capital cost for subsequent periods. To calculate the risk margin $RM_{0}$, we insert (36) into (39) to derive a formula for $SCR_{0}$. This can then be inserted into (36) and solved for $RM_{0}$ to give:
$RM_{0}=\frac{((1+r_{0,1}+\eta)-P[M_{1}]\cdot(1+r_{0,1}))\cdot\rho(A_{1}^{BEL}-L_{1})}{(1+r_{0,1}+\eta)\cdot(1+r_{0,1})}+\frac{\mathbb{E}[1_{M_{1}}\cdot RM_{1}]}{(1+r_{0,1}+\eta)}$ (40)
Third stage. We make the further simplification:
  1. (f)
    The financiability condition with equality is replaced (by leaving out $(\cdot)_{+}$) by the stricter condition $\mathbb{E}[AC_{1}^{pre}]=(1+r_{0,1}+\eta)\cdot SCR_{0}$.
This gives a lower required capital $SCR_{0}$ and thus a higher production cost than without the simplification and in this sense an upper bound. Formula (39) then simplifies (as $M_{1}$ disappears), and solving it for the risk margin $RM_{0}$:
@@
So the risk margin is the sum of the discounted capital cost for the period from 00 to $1$ and the discounted expected capital cost for the periods after $1$, which is a first step for deriving formula (33) by suitable recursion. As above, $RM_{0}$ cannot directly be calculated from this but from (40) (with $M_{1}$ disappearing):
$RM_{0}=\frac{\eta\cdot\rho(A_{1}^{BEL}-L_{1})}{(1+r_{0,1}+\eta)\cdot(1+r_{0,1})}+\frac{\mathbb{E}[RM_{1}]}{(1+r_{0,1}+\eta)}$ (42)
Further assuming that also the assets for the best estimate are invested in the risk-free one-year zero-coupon bond, i.e. $A_{1}^{BEL}=(1+r_{0,1})\cdot BEL_{0}$, the correspondingly simplified condition (38) directly gives